OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim? If D = dad's age now and J = Jim's age now, which of the following systems could be used to solve the problem? D = 4J and D + 10 = 2J + 20 D = 4J and D + 10 = 2(4J + 10) + 20 D = 4J and D + 10 = 2(J + 10) + 20

OpenStudy (anonymous):

Last choice is 10 years dads age is D+10 and Jimmy's will be (J+10) and since dad will be twice his age so the equ is D+10=2(J+10)

OpenStudy (anonymous):

And since he'll be 20 years older than the twice so we will get D+10=2(J+10)+20

OpenStudy (anonymous):

in*